OpenStudy (anonymous):

The equation of a circle is shown below. x2 + (y - 9)2 = 4 Which statement is true about the center of the circle? * It is located on the x-axis. * It is located on the y-axis. * It is located in the third quadrant. * It is located in the fourth quadrant.

OpenStudy (anonymous):

looks like \[(x-h)^2+(y-k)^2=r^2\] a circle with center \((h,k)\) and radius \(r\) write it as \[(x-0)^2+(y-9)^2=2^2\] to identify the center as \((0,9)\) and the radius as 2

OpenStudy (anonymous):

since the first coordinate is \(x=0\) this is located on the \(y\) axis

OpenStudy (anonymous):

thanks (:, is it possible that for my knowledge for future questions you tell me where are the quadrants located because i really don't know.